OpenStudy (anonymous):

Simplify: (3a^3b^4)2(b^3)^3

OpenStudy (anonymous):

(3a^3b^4)2(b^3)^3 =(3a^3b^4)2(b^9) = 6a^3b^4)(b^9) =6a^3b^13

OpenStudy (anonymous):

\[(3a^3b^4)2(b^3)^3=6a^3b^4b^{3\cdot3}=6a^3b^{4+9}=6a^3b^{13}\]